# Quantum circuits with free fermions in disguise

### v2

@@ 1 Introduction
The structure of the paper is as follows. In Section 2 we discuss the free fermionic properties in general, and we formulate our main questions. In Section 3 we consider Jordan-Wigner diagonalizable models; here we just collect well known results. In Section 4 we introduce the FFD algebra (Free Fermions in Disguise), on which the 4-fermion model of Fendley was built. We discuss basic properties of this algebra, together with new results for the “minimal” representations. We also collect the key results of [6] about the solutions of these models. In Section 5 we present a free fermionic quantum circuit within the FFD algebra, derived from the results of [6]. In Section 6 we investigate the free fermionic circuits for small system sizes. Finally, in Section 7 we present new free fermionic circuits with different geometries; here the free fermionic nature is supported by numerical data. In Section 8 we treat the model of [17]: we factorize the transfer matrix, thereby we compute a unitary quantum circuit with local gates also for this model. Finally, our conclusions are presented in Section 9, and a few details about the construction of the free fermionic operators are provided in Appendix A. Appendix B contains the numerical data supporting the conjecture in Section 7. Appendix C provides a detailed proof of the factorization of the transfer matrix for the generalized model [17].
@@ 2.3 Posing the problem
$u_{j}=e^{i\varphi_{j}h_{j}}=\cos(\varphi_{j})+i\sin(\varphi_{j})h_{j}.$ (2.8)
We then construct various quantum circuits as products of these local gates. Let $\mathcal{S}=\{s_{1},s_{2},\dots,s_{N}\}$ be a string of indices, with length $N$, such that $s_{k}\in[1,M]$. Repetitions of indices are allowed, and generally we expect that $N$ will grow linearly with $M$. Then we take the product
@@
Let us assume that each index $k$ appears in $\mathcal{S}$ the same number of times (the number of repetitions is given by $\rho=M/N$), and let us substitute
@@
Below we will see that there exist free fermionic unitary circuits for which the lowest order commuting charge is actually a Hamiltonian of the form (2.7). In such a case there is a direct mapping between the set of angles $\{\varphi_{j}\}$ and the parameters $\{\alpha_{j}\}$ of the Hamiltonian. However, generally it is not possible to find a Hamiltonian of the form (2.7) commuting with $\mathcal{V}$, and the “smallest” commuting charge emerges at a slightly higher order.
@@ 2.4 Basic properties of the circuits
$u_{1}u_{1}=e^{i\varphi_{1}h_{1}}e^{i\varphi_{1}h_{1}}=e^{i2\varphi_{1}h_{1}}.$ (2.16)
@@
$u_{1}u_{2}u_{3}u_{2}u_{1}$ (2.17)
for some triplet of operators $h_{1},h_{2},h_{3}$, such that $[h_{3},h_{2}]=0$. Then by Property 1. and 2. we are free to investigate the circuit $u_{3}u_{2}u_{2}u_{1}u_{1}$, and by Property 3. we can simplify it and investigate $u_{3}u_{2}u_{1}$ instead.
@@ 4.2 Solution of the Hamiltonians in the FFD algebra
where we introduced
@@ 5 FFD algebra: a circuit from the transfer matrix
$P_{M}(v^{2})=P_{M-1}(v^{2})-v^{2}\frac{\sin^{2}(2\varphi)}{\cos^{6}(2\varphi)}P_{M-3}(v^{2}),\quad M\geq 3$ (5.17)
@@ 8 A free fermionic circuit in the generalized model
The theoretical relevance of the model lies in the fact that the Hamiltonian does not directly fit into the framework of [12, 13]. There are two differences. First, the frustration graph of the model does not satisfy the conditions given in [12, 13] for free fermion solvability. Second, the pre-factors for the different terms in the Hamiltonian are not independent from each other. In fact, for a given $L$ there are $23L-$ 2 terms in the Hamiltonian, but only $2L$ coupling constants. It was remarked in [12, 13] that their conditions are sufficient but not necessary for the free fermionic solvability, and that fine tuned cases do exist which are free fermionic only if there are special relations between the pre-factors in the Hamiltonian. The model of [17] presents one such case.
@@ 8.1 Transfer matrix and free fermions in the generalized model
$\Psi_{\pm k}=\frac{1}{N_{k}}T(\mp\tilde{v}_{k})\alpha_{L+1}T(\pm\tilde{v}_{k})$ (8.12)
@@ 9 Conclusions
In this work we studied unitary quantum circuits with a hidden free fermionic spectrum. Our central result is that a quantum circuit with the specific structure $G\cdot G^{T}$ is free fermionic in many regular geometries, both in the original 4-fermion model of Fendley, and in the generalized model of [17]. In the case of the linear depth circuits we proved the free fermionic property analytically, while for the constant depth circuits it was demonstrated numerically.
@@
We leave these questions to further work.

### v3

@@ 1 Introduction
The structure of the paper is as follows. In Section 2 we discuss the free fermionic properties in general, and we formulate our main questions. In Section 3 we consider Jordan-Wigner diagonalizable models; here we just collect well known results. In Section 4 we introduce the FFD algebra (Free Fermions in Disguise), on which the 4-fermion model of Fendley was built. We discuss basic properties of this algebra, together with new results for the “minimal” representations. We also collect the key results of [6] about the solutions of these models. In Section 5 we present a free fermionic quantum circuit within the FFD algebra, derived from the results of [6]. In Section 6 we investigate the free fermionic circuits for small system sizes. In Section 7 we present new free fermionic circuits with different geometries; here the free fermionic nature is supported by numerical data. Finally, in Section 8 we treat the generalization of the FFD algebra [17] and the free-fermionic model built on that algebra. We factorize the transfer matrix obtained in [17], i.e., we represent the transfer matrix as a product of local gates in the generalized FFD algebra. Then we compute a unitary quantum circuit also for this model. Finally, our conclusions are presented in Section 9, and a few details about the construction of the free fermionic operators are provided in Appendix A. Appendix B contains the numerical data supporting the conjecture in Section 7. Appendix C provides a detailed proof of the factorization of the transfer matrix for the generalized model [17].
For the convenience of the reader we summarize here the most important new results of this work:
  • In Section 5 we derive a unitary staircase circuit for the FFD model, building on the results of [6].
  • In Section 7 we obtain new free fermionic quantum circuits in the FFD algebra, including a circuit with finite depth.
  • In Section 8, we derived the factorized form of the transfer matrix obtained in [17], from which we obtain a free fermionic staircase circuit in the generalized FFD algebra.
@@ 2.3 Posing the problem
$u_{j}({\varphi_{j}})=e^{i\varphi_{j}h_{j}}=\cos(\varphi_{j})+i\sin(\varphi_{j})h_{j}.$ (2.8)
We note that the dependence of $\varphi_{j}$ on $u_{j}(\varphi_{j})$ is often omitted in the following for simplicity. We then construct various quantum circuits as products of these local gates. Let $\mathcal{S}=\{s_{1},s_{2},\dots,s_{N}\}$ be a string of indices, with length $N$, such that $s_{k}\in[1,M]$. Repetitions of indices are allowed, and generally we expect that $N$ will grow linearly with $M$. Then we take the product
@@
Let us assume that each index $k$ appears in $\mathcal{S}$ the same number of times (the number of repetitions is given by $\rho=N/M$), and let us substitute
@@
Below we will see that there exist free fermionic unitary circuits for which the lowest order commuting charge is actually a Hamiltonian of the form (2.7). In such a case there is a direct mapping between the set of angles $\{\varphi_{j}\}$ and the parameters $\{\alpha_{j}\}$ of the Hamiltonian, which reduces to (2.13) in the infinitesimal angle limit. However, generally it is not possible to find a Hamiltonian of the form (2.7) commuting with $\mathcal{V}$, and the “smallest” commuting charge emerges at a slightly higher order.
@@ 2.4 Basic properties of the circuits
$u_{1}{(\varphi_{1})}u_{1}{(\varphi_{1})}$ $=e^{i\varphi_{1}h_{1}}e^{i\varphi_{1}h_{1}}=e^{i2\varphi_{1}h_{1}}$
$={u_{1}(2\varphi_{1})}.$ (2.16)
@@
${u_{1}(\varphi_{1})u_{2}(\varphi_{2})u_{3}(\varphi_{3})u_{2}(\varphi_{2})u_{1}(\varphi_{1})}$ (2.17)
for some triplet of operators $h_{1},h_{2},h_{3}$, such that $[h_{3},h_{2}]=0$. Then by Property 1. and 2. we are free to investigate the circuit $u_{3}(\varphi_{3})u_{2}(\varphi_{2})u_{2}(\varphi_{2})u_{1}(\varphi_{1})u_{1}(\varphi_{1})$, and by Property 3. we can simplify it and investigate $u_{3}(\varphi_{3})u_{2}(2\varphi_{2})u_{1}(2\varphi_{1})=u_{3}(\varphi^{\prime}_{3})u_{2}(\varphi^{\prime}_{2})u_{1}(\varphi^{\prime}_{1})$ instead.
@@ 4.2 Solution of the Hamiltonians in the FFD algebra
where we introduced the quasi-particle energies $\varepsilon_{k}$ as
@@ 5 FFD algebra: a circuit from the transfer matrix
$P_{M}(v^{2})=P_{M-1}(v^{2})-v^{2}\frac{\sin^{2}(2\varphi)}{\cos^{6}(2\varphi)}P_{M-3}(v^{2}),\quad M\geq 3.$ (5.17)
@@ 8 A free fermionic circuit in the generalized model
This model is important for the purely theoretical questions of free fermion solvability. The relevance of the model lies in the fact that the Hamiltonian does not directly fit into the framework of [12, 13]. There are two differences. First, the frustration graph of the model does not satisfy the conditions given in [12, 13] for free fermion solvability. Second, the pre-factors for the different terms in the Hamiltonian are not independent from each other. In fact, for a given $L$ there are $23L-$ 2 terms in the Hamiltonian, but only $2L$ coupling constants. It was remarked in [12, 13] that their conditions are sufficient but not necessary for the free fermionic solvability, and that fine tuned cases do exist which are free fermionic only if there are special relations between the pre-factors in the Hamiltonian. The model of [17] presents one such case.
@@ 8.1 Transfer matrix and free fermions in the generalized model
$\Psi_{\pm k}=\frac{1}{N_{k}}T(\mp\tilde{v}_{k})\alpha_{L+1}T(\pm\tilde{v}_{k}).$ (8.12)
@@ 9 Conclusions
In this work we studied unitary quantum circuits with a hidden free fermionic spectrum. Our central result is that a quantum circuit with the specific structure $G\cdot G^{T}$ is free fermionic in many regular geometries, both in the original 4-fermion model of Fendley, and in the generalized model of [17]. In the case of the linear depth circuits we proved the free fermionic property analytically, while for the constant depth circuits it was demonstrated numerically. Our results could serve as a starting point for the study of non-equilibrium dynamics generated by the free fermionic circuits.
@@
A further interesting question is whether there can be quantum circuits with hidden free fermions, such that the circuits are still built from local gates, but there is no local Hamiltonian associated with the circuit. We can not rule out such possibilites, but our current methods are adapted to those cases when there is a local Hamiltonian in the model, at least for the infinitesimal angle limit. Alternatively, it might also happen to certain circuit geometries become free fermionic for special choices of the local rotation angles. Furthermore, it is a nontrivial question whether there still exists a local Hamiltonian that commutes with the quantum circuit even when the angles are special finite numbers. Our treatment does not cover such cases.
We leave the aforementioned open questions to further work.